OpenStudy (dls):

An urn contains 5 red and 5 black balls.A ball is drawn at random,its colour is noted and is returned to the urn.Moreover,2 additional balls of the colour drawn are put in the urn and then a ball is drawn at random.What is the probability that the 2nd ball is red?

ganeshie8 (ganeshie8):

P(if first ball is RED case) + P(if first ball is BLACK case)

ganeshie8 (ganeshie8):

P(if first ball is RED case) : (5/10)(7/12)

OpenStudy (dls):

Case 1: Let Red ball is drawn.. 5/10=1/2 After adding 2 balls..7/12 ------------------------------ Case 2: Let black ball is drawn.. 5/10=1/2 After adding 2 balls..7/12 yet again..

ganeshie8 (ganeshie8):

P(if first ball is BLACK case) : (5/10)(5/12)

OpenStudy (dls):

:O

OpenStudy (dls):

in 2nd case..P(red ball)=5/12

ganeshie8 (ganeshie8):

in case 2, you're not adding Red ball. so after adding 2 black balls, probability wud be 5/12

ganeshie8 (ganeshie8):

yeah

OpenStudy (dls):

okay thanks :D i have another similar question

ganeshie8 (ganeshie8):

il try, dint do much probability recently..

OpenStudy (dls):

Suppose a girl throws a die.If she get s 5 or 6 she tosses a coin three times and notes the no. of heads.If she gets 1,2,3 or 4 she tosses a coin once and notes whether a head or tail is obtained..If she obtained exactly 1 head what is the prob that she threw 1,2,3 or 4 with the die?

ganeshie8 (ganeshie8):

P(coins) thing is given = 1 head

ganeshie8 (ganeshie8):

i mean the outcome was known upfront for that event

ganeshie8 (ganeshie8):

looks hard to meh